OpenStudy (anonymous):

write a fraction equivalent to 11/24 ?

OpenStudy (anonymous):

220/480

OpenStudy (anonymous):

can i have a medal

OpenStudy (anonymous):

bruh

OpenStudy (minamoe):

what grade are you in?

OpenStudy (amorfide):

lmao the format messed up there

OpenStudy (amorfide):

to get a fraction of the same value you must multiply the fraction by a fraction with the same number in the numerator and denominator for example you have \[\frac{ 11 }{ 24 }\] multiply by a fraction with the same number top and bottom for example 2/2 or 3/3 \[\frac{ 11 }{ 24 }\times \frac{ 2 }{2 }\]

OpenStudy (amorfide):

this gives \[\frac{ 22 }{ 48}\]

OpenStudy (amorfide):

because the fraction we multiplied by is also equal to 1 that means we didnt change the value of the original fraction meaning 11/24 = 22/48

OpenStudy (elizadaniel):

33/72 , 44/96, 22/48